OpenStudy (anonymous):

OKSY HOW DO I FIND (X,Y,Z).... 3X+4Y+2Z=24

OpenStudy (anonymous):

AND (X+3)^2+Z^5+25

OpenStudy (anonymous):

are they both part of the same system?

OpenStudy (anonymous):

YEA THEY ARE

OpenStudy (anonymous):

WELL WHAT I WAS DOING WAS SOLVE INDIVIDUALY LIKE 3X=24 AND THEN DIVIDING IT TO BOTH SIDES WICH GAVE ME 8 SO X=8 AND THEN I DID THAT TO Y AND Z

OpenStudy (anonymous):

IM NOT SURE IF THATS CORRECT

OpenStudy (anonymous):

No. What is your second equation up there equal to?

OpenStudy (anonymous):

I GOT SOMETHING LIKE (-3,0,5)

OpenStudy (anonymous):

AND FOR MY FRST EQUATION I GOT (8,6,12)

OpenStudy (anonymous):

No. Those answers aren't correct. If you can tell me what your second equation is equal to then I can help you.

OpenStudy (anonymous):

OOOH SRRY THE SEC ONE EQUALS 25 I ACCENDENTLY PUT PLUS

OpenStudy (anonymous):

No problem. Solve your first equation for x... x=8-(4/3)y-(2/3)z...and so... x+3=11-(4/3)y-(2/3)z... do you follow?

OpenStudy (anonymous):

yea a little

OpenStudy (anonymous):

well, this gets kind of messy, so it's important that you can get that far, because you also solve it for z & substitute each into your second equation.

OpenStudy (anonymous):

HEHE OKAY THNX SO MUCH I JUS HAD TO LOOK AT IT A COUPLE OF TIMES TO UNDERSTAND IT THANK YU